The times t required to do a job varies inversely as the number of people P working. It takes 6 hours for 9 volunteers to pick up trash from 1 mile of roadway. How long would it take 10 volunteers to complete this job?
Answer by stanbon(75887)

The times t required to do a job varies inversely as the number of people P working.
t = k/p
------
Solve for "k" using "It takes 6 hours for 9 volunteers to pick up trash from 1 mile of roadway".
6 = k/9
k = 54
--------------------
Equation::
t = 54/p
---------------------
How long would it take 10 volunteers to complete this job?
t = 54/10 = 5.4 days
-----------------------
Cheers,
Stan H.
-----------------------